1. Triumph Corporation extends credit to many of its customers, and sets the term of these credits at:

1.45/7; net/29

Jones is a customer of Triumph and has received the above credit on a recent purchase of $100,000. Jones is trying decide whether to take Triumph up on its offer to pay early or instead to wait until the last possible day to pay the amount in full.

Help Jones out by computing the annual interest rate implied by these credit terms.
%

Place your answer as a percentage without any percent sign. That is, if you think the answer is twenty five point five percent, then place your answer as 25.5

2.Crisps has received an order for 13500 bags of potato chips from BigBag. Crisps views BigBag to be a one-time customer. Crisps sells its large bags of potato chips for $2.15 each, and calculates its internal cost for the product at $1.05 each.

Market research estimates that there is a 37% chance that BigBag will pay in full what it owes.
Based on this information, what is the NPV to Crisps of offering credit to BigBag?
$

Place your answer to the nearest dollar. Do not include a dollar sign or comma.

3.Crisps has received an order for 16500 bags of potato chips from BigBag. Crisps views BigBag to be a long-term customer and believes they will continue to place the same order year after year forever. Crisps sells its large bags of potato chips for $2.25 each, and calculates its internal cost for the product at $1.05 each.

Market research estimates that there is a 27% chance that BigBag will pay in full what it owes. Crisps uses a discount rate of 6.85% for all NPV analysis.

Based on this information, calculate the NPV of this credit decision?
$

Place your answer to the nearest dollar. Do not use a Dollar sign or commas within your answer.
Answer #1

Answer:

Question 1)

Given Discount rate r= 1.45 % if payment is made in 7 days

Discount days D= 7 days

Allowable Payment days N= 29 days

Annual interest rate= r*360/{(1-r)*(N-D)}=1.45%*360/{(1-1.45%)*(29-7)}=24.08